Fermi National Laboratory
 

Accelerator Update for June 11, 2007 - June 13, 2007

 

Monday June 11

The day shift began with Operations monitoring the Recycler stash, with the Antiproton Source (Pbar) stacking, with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) waiting for the colliding beams.

Operations put the Tevatron (TeV) into shot setup at 8:14 AM.

Operations established store 5485 at 10:14 AM with an initial luminosity of 236.84E30.

At 10:24 AM, a MI kicker (KPS1A) tripped off and caused the MI beam permit to trip. Everything reset.

A Pbar expert began an Accumulator orbit study at 11:26 AM.

At 12:23 PM, a Linac RF station (LRF2) tripped off. Operators and experts investigated and had the station back on about 30 minutes later.

At 12:47 PM, Operations reported that store 5485 was lost due to a vacuum problem at TeV sector A0.

At 12:51 PM, the cryo system technicians at CHL reported that they lost the sector A0 compressors. They brought on the spare compressors at B0 to compensate.

At 1:01 PM, CHL reported that the electrical yard breaker at sector A0 was off. Operations contacted the Duty Electrician. This turned out to be the first cause for the loss of store 5485.

Operations had CDF and D0 prepared for an access by 1:39 PM.

Operations reported at 1:56 PM that the Duty Electrician managed, on the second attempt, to reset the TeV sector A0 yard breaker.

At 3:44 PM, LRF2 tripped off. Operators and experts investigated.

CDF came out of access at 3:47 PM and D0 came out a few minutes later.

LRF2 came back on line and Linac resumed sending beam at 4:14 PM.

Operations reported at 4:16 PM that the Recycler had stashed antiprotons twice during the day shift.

At 4:31 PM, the MI-10 kicker tripped off. It reset.

Operations put the TeV through a dry squeeze at 4:43 PM.

TeV experts began a study period at 5:40 PM.

At 6:17 PM, the MI-40 abort kickers tripped off and took down the MI beam permit. Everything reset.

At 6:37 PM, a quadrupole power supply (I:Q703) in the P1 transfer line tripped off and caused the NuMI beam permit to trip off. Everything reset.

TeV experts completed their study at 7:16 PM. Operations put the TeV through a dry squeeze.

Operations put the TeV into shot setup at 7:48 PM.

At 8:23 PM, a MI RF station (MIRF13) tripped off and would not reset. Operators bypassed the station and turned on MIRF17 to take its place.

At 8:59 PM, the MI power supply interlock loop tripped off. It reset.

Operations established store 5488 at 9:58 PM with an initial luminosity of 132.35E30.

Operations reported at 11:59 PM that the Recycler had stashed three times during the evening shift.

Tuesday June 12

The midnight shift began with Operations monitoring store 5488 (94.27E30) and the stash (41.2E10), with Pbar stacking (42.0E10), and with NuMI and MiniBooNE taking beam.

At 12:48 AM, the MI beam permit tripped off for the first of two times in seven minutes due to kicker trips (KPS4A & KPS4B). Everything reset.

At 2:05 AM, KPS1A tripped off. It reset.

At 2:53 AM, KPS1A & B tripped off. They reset.

At 3:06 AM, a Tevatron Electron Lens (TEL2) tripped off due to a vacuum interlock. It reset.

At 3:09 AM, a TeV RF station (TRF8) tripped off. RF experts will investigate in between stores.

At 3:33 AM, KPS1A tripped off. It reset.

At 4:58 AM, KPS1A tripped off. It reset.

At 5:23 AM, KPS4A & B tripped off. They reset.

At 6:37 AM, KPS1A tripped off. It reset.

At 7:24 AM, a Pbar Debuncher quadrupole power supply tripped off and would not reset. Operators investigated. An expert came in and had the station back on 30 minutes later.

Operations reported at 7:56 AM that the Recycler had stashed three times during the midnight shift.

An RF expert brought MIRF13 back on line at 8:20 AM. MIRF17 was put back in reserve.

At 8:48 AM, KPS1A tripped off. It reset.

At 10:40 AM, KPS1A tripped off. It reset.

Operations terminated store 5488 at 12:05 PM.

Operations put the TeV into shot setup at 12:20 PM.

At 12:29 PM, I:Q703 tripped off and caused the P1 transfer line to trip as well as causing the NuMI beam permit to trip. Everything reset.

Operations prepared the Pelletron for access at 2:22 PM.

Operations established store 5489 at 2:33 PM with an initial luminosity of 177.62E30.

An MI expert began a slipstacking study at 2:59 PM.

Recycler experts began a linear bucket study at 3:10 PM.

Operations reported at 4:15 PM that the Recycler had stashed two times during the day shift.

The Recycler studies ended at 4:33 PM.

At 8:10 PM, the Linac Buncher tripped off for the first of two times in about 30 minutes. It reset.

At 9:36 PM, KPS1A tripped off. It reset.

Operations reported at 11:59 PM that the Recycler had stashed four times during the evening shift.

Wednesday June 13

The midnight shift began with Operations monitoring store 5489 (64.42E30) and the stash (196.0E10), with Pbar stacking (17.7E10), and with NuMI and MiniBooNE taking beam.

At 1:01 AM, KPS1A tripped off. It reset.

At 4:56 AM, Operations reported that the Recycler had lost 1E10 of its stash during a transfer.

At 5:04 AM, KPS1A tripped off. It reset.

At 5:10 AM, KPS1B tripped off. It reset.

At 6:40 AM, Operations reported that the Recycler had lost 1E10 of its stash during a transfer.

At 6:55 AM, a Booster vacuum pump (B:IPRF17) tripped off and took down BRF7 & 8. Everything reset.

Operations reported at 7:55 AM that the Recycler had stashed three times during the midnight shift.

The Plans for Wednesday

The plans for today are to stack, stash, and store. The Run Coordinator will allow a MI study. There is a planned shutdown tomorrow, 6/14/07, to find and repair the MI RF LCW leak.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/13/2007    email Fermilab